# MQED-QD: An Open-Source Package for Quantum Dynamics Simulation in Complex Dielectric Environments

### 05 Mar 2026 v1

@@ 5 MQED-QD Package Details
The representative command-line entry points are listed in Table 1. The MQED-QD package is openly accessible on our Github repositoryGitHub-repo and supports the standard Python package installer (PIP).
@@ 6.1 Simulation Details
To demonstrate the robustness of the workflow, we consider a chain of molecular emitters in close proximity to a silver nanostructure. To be consistent with the experimental reports and common practice in the MQED literature, we evaluate the dyadic Green’s functions in SI units using either the analytical solution or the calibrated BEM calculations. The dielectric function of silver nanostructure is given by the MNPBEM toolbox.34 The MQED-QD package allows users to define the dielectric functions as an input file dielectric.xlxs.
@@ 7 Conclusion
We look forward to expanding the functionalities of the MQED-QD package to address a broader range of complex physical phenomena. Specifically, we aim to develop the following features. (i) The integration of wavefunction-based methods11 or few-mode models54, 10 is essential for capturing non-Markovian quantum dynamics, which is critical for accurately simulating strongly coupled plasmon–polariton systems. (ii) Incorporating nuclear degrees of freedom via the hierarchical equations of motion (HEOM) or tensor-network-based methods will enable the investigation of vibronic coupling and non-adiabatic effects at the light–matter interface. (iii) Developing more efficient machine-learning-based surrogate models will facilitate the bypass of direct dyadic Green’s function constructions. We expect these developments will bridge the gap between the nanophotonics and quantum dynamics communities.

### 06 Mar 2026 v2

@@ 5 MQED-QD Package Details
The representative command-line entry points are listed in Table 1. The MQED-QD package is openly accessible on our Github repository1 and supports the standard Python package installer (PIP).
@@ 6.1 Simulation Details
To demonstrate the robustness of the workflow, we consider a chain of molecular emitters in close proximity to a silver nanostructure. To be consistent with the experimental reports and common practice in the MQED literature, we evaluate the dyadic Green’s functions in SI units using either the analytical solution or the calibrated BEM calculations. The dielectric function of silver nanostructure is given by the MNPBEM toolbox.35 The MQED-QD package allows users to define the dielectric functions as an input file dielectric.xlxs.
@@ 7 Conclusion
We look forward to expanding the functionalities of the MQED-QD package to address a broader range of complex physical phenomena. Specifically, we aim to develop the following features. (i) The integration of wavefunction-based methods12 or few-mode models55, 11 is essential for capturing non-Markovian quantum dynamics, which is critical for accurately simulating strongly coupled plasmon–polariton systems. (ii) Incorporating nuclear degrees of freedom via the hierarchical equations of motion (HEOM) or tensor-network-based methods will enable the investigation of vibronic coupling and non-adiabatic effects at the light–matter interface. (iii) Developing more efficient machine-learning-based surrogate models will facilitate the bypass of direct dyadic Green’s function constructions. We expect these developments will bridge the gap between the nanophotonics and quantum dynamics communities.
@@
The source code of the MQED-QD package is openly available on our Github repository: https://github.com/MQED-QD/Macroscopic-Quantum-Electrodynamics